# Transverse Momentum-Dependent Heavy-Quark Fragmentation at Next-to-Leading Order

### v1

@@ 4.1 All-order renormalization properties
where $S^{(n_{\ell})}$ is the bare TMD soft function in the theory where only the $n_{\ell}$ light quark degrees of freedom are active. This is the same soft function as appears in light-quark TMD PDFs or FFs at transverse momenta below the heavy-quark threshold: The fact that it also arises here and cancels the rapidity divergences as $\eta\to 0$ in the bare light-spin density matrix follows from consistency of the TMD factorization for flavor-changing hard processes, e.g. charged-current DIS $e^{-}d\to\nu_{e}\,c$111 We assume as in section 2.1 that a rapidity regulator $\eta$ is used for which soft (zero-bin) subtractions in the bare collinear matrix element are scaleless. If this is not the case, it is crucial that zero-bin subtractions are defined by modes that are soft in a frame where the heavy quark is highly boosted, i.e., their interactions with the heavy-quark sector are encoded in lightlike Wilson lines and the multiplicative zero-bin subtraction factor is obtained by taking $\bar{n}\cdot v\to\infty$. The UV $\overline{\text{MS}}$ renormalization factor $Z_{\chi_{1}}$ likewise follows from consistency and is given by
@@
A natural boundary condition for $\rho_{\ell}$, where it is free of large rapidity logarithms, is given by $\rho\sim 1$, i.e., the rest frame of the heavy quark. By contrast, in physical cross sections, the boost is fixed to $\rho=Q/m\gg 1$, i.e., the lab frame in the $e^{+}e^{-}$ case, or the Breit frame for SIDIS. The $\rho$-evolution between those scales resums the associated large rapidity logarithms. As in the relativistic case, the two-dimensional evolution equations satisfy an integrability condition,
@@ 4.3 One-loop calculation
Figure -1076: Left: Diagrams contributing to the total bHQET TMD fragmentation factor in the perturbative domain. Wilson lines along the lightlike ${\bar{n}}$ direction are indicated by the slanted double lines. The horizontal double lines indicate the Wilson lines along the timelike $v$ direction. The dashed line indicates on-shell cuts after introducing a complete set of states. The mirror diagram for (a) is understood and included in expressions given in the text. Right: Phase-space integrands $\mathcal{A}$ for the two real-emissions diagrams as a function of the gluon emission rapidity $y$.
@@
in analogy to eq. (3.10). Virtual diagrams with $X$ the vacuum state are scaleless. The two contributing real emission diagrams at $\mathcal{O}(\alpha_{s})$ in Feynman gauge are shown in figure -1076b. The one-gluon Feynman rule for the $Y_{v}$ Wilson line can e.g. be read off from ref. [42], which uses a covariant derivative convention consistent with the one-gluon Feynman rule we use for $W_{\eta}$. Including the mirror diagram for diagram (a), the two diagrams are both given by simple integrals over the rapidity $y$ of the emitted gluon in the heavy-quark rest frame,
@@
The origin of the rapidity divergence, like the structure of the integrands as a whole, has a rather intuitive interpretation, which we illustrate in the right panel of figure -1076b: Emissions at large positive gluon rapidity $y$ in the rest frame of the heavy quark are strongly suppressed for either contributing diagram, which is the well-known dead-cone effect [28]. For diagram (b), where the static heavy quark $Y_{v}$ both emits and absorbs the gluon, the contribution to the emission probability (dot-dashed blue) peaks at $y=0$ and is suppressed also at large negative rapidities because the contribution must be even in the rest frame. In contrast to this, the contribution from diagram (a), where the emission couples to the $W_{\eta}$ that arises from integrating out the interactions with the anticollinear and soft sectors and is required for gauge invariance, tends to a large constant as $y\to-\infty$ (dashed red), leading to a rapidity divergence. Of course, this precisely corresponds to the central soft momentum region, and thus after introducing the regulator $e^{-\eta y}$ with $\eta<0$ to dampen the behavior as $y\to\infty$ (solid red), the associated pole can cancel against the soft function as usual.
@@ 6 Joint threshold and $q_{T}$ resummation for massive quark fragmentation
Figure -1075: Relevant modes contributing to the heavy-quark TMD FF as the collinear momentum fraction $z_{H}\to 1$ while counting $m\sim k_{T}$. See the text for a detailed discussion.
@@
We stress that we continue to count $m\sim k_{T}\sim 1/b_{T}$. In figure -1075 we illustrate the contributing modes in the EFT for this scenario. Figure -1075 (a) shows the original, unmodified mode setup that led to eq. (2.1). Here the partonic heavy-quark TMD FF arises from matching collinear modes (green), accounting for the usual contribution of the central soft modes (orange), onto the nonperturbative modes (gray) describing the brown muck, which is isotropic in the rest frame of the heavy quark indicated by the dashed black line. The collinear modes are kinematically unconstrained in the longitudinal direction because $\bar{z}_{H}\sim 1$.
Once the modified assumption eq. (6.1) is made, collinear modes at the scale $\mu\sim k_{T}\sim m$ can no longer be radiated into the final state. Their effect becomes purely virtual and is encoded in the matching coefficient $C_{m}$ (green cross) from matching onto bHQET modes at yet lower virtuality. Real emissions into the $n$-collinear direction at the scale $\mu\sim k_{T}\sim m$ can instead only have collinear-soft (pink) scaling
@@
which is the characteristic mode of double-differential $z\to 1$ threshold and $q_{T}$ resummation, often simply dubbed joint resummation [60, 61, 62, 63, 64, 65, 66, 67, 68].222 We note that while ref. [64] also dealt with final-state heavy quarks, that work considered the joint resummation for the total transverse momentum of a heavy quark pair with respect to the beam direction, which is a problem that is distinct from the one we consider here.
The physical picture at scales $\mu\ll k_{T}\sim m$ in this scenario, which is governed by one or several bHQET modes at their characteristic lab-frame rapidity $p^{-}/p^{+}\sim Q^{2}/m$, further depends on the relative size of $\Lambda_{\mathrm{QCD}}/m$ and $\bar{z}_{H}$, specifically whether $\Lambda_{\mathrm{QCD}}/m\sim\bar{z}_{H}$ or $\Lambda_{\mathrm{QCD}}/m\ll\bar{z}_{H}$. (Since $\Lambda_{\mathrm{QCD}}$ introduces a fourth scale into the problem, i.e., a third independent dimensionless power counting parameter, either case may hold in general and independent of $k_{T}/Q$ and $m/Q$.) The relevant modes for the case $\Lambda_{\mathrm{QCD}}/m\sim\bar{z}_{H}$ are illustrated in figure -1075 (c), and are in fact closely analogous to the well-understood case of single-inclusive heavy-hadron production near the endpoint in $e^{+}e^{-}$ collisions. Specifically, the nonperturbative fragmentation process at the scale $\mu\sim\Lambda_{\mathrm{QCD}}\sim\bar{z}_{H}m$ provides a leading contribution to the measured longitudinal hadron momentum in this case (but not to its transverse momentum, where the contribution is power suppressed!). This effect is encoded in the well-known collinear bHQET fragmentation shape function $F_{H}(k^{-},\mu)$ [5, 6]. [^3] We use the symbol $F_{H}$ here to distinguish the shape function from the TMD soft function; in ref. [6], the symbol $S_{H/Q}$ is used instead, and our partonic $\hat{F}$ corresponds to their $S_{Q/Q}$. In total, the factorization of the hadron-level heavy-quark TMD FF in this regime reads, with $\zeta=\omega^{2}$ the Collins-Soper scale,
$D_{1\,H/Q}(z_{H},b_{T},\mu,\omega^{2})$ $=C_{m}(m,\mu,\omega^{2})\,\int\!\mathrm{d}\ell^{-}\,\mathcal{S}\Bigl{(}b_{T},m,\mu,\frac{\ell^{-}}{\nu}\Bigr{)}\,\sqrt{S}(b_{T},m,\mu,\nu)$
×$\quad\times\int\!\frac{\mathrm{d}k^{-}}{\bar{n}\cdot v}\,F_{H}\Bigl{(}\frac{k^{-}}{\bar{n}\cdot v},\mu\Bigr{)}\,\delta\Bigl{(}\bar{z}_{H}-\frac{\ell^{-}+k^{-}}{\omega}\Bigr{)}$
@@
where $\mathcal{S}$ and $\sqrt{S}$ are the separately renormalized collinear-soft function from joint resummation [68] and the square root of the standard TMD soft function, respectively. By evolving all terms in eq. (6) from their canonical scales (e.g. in Mellin space $N\leftrightarrow z_{H}$ where $\nu_{\mathcal{S}}\sim\bar{n}\cdot v\,\mu_{F_{H}}\sim Q/N$ and the evolution is multiplicative) to the common scale $\mu\sim\sqrt{\zeta}\sim Q$ with the hard function, all large threshold and transverse momentum logarithms present in the heavy-hadron TMD cross section in this regime can be resummed simultaneously.
It is instructive to compare the above to the refactorization of the light-parton TMD FF in the joint resummation regime $k_{T}\ll\bar{z}_{h}Q\ll Q$, which follows from crossing the light-parton TMD PDF refactorization relations [68],
$D_{1\,h/q}(z_{h},b_{T},\mu,\omega^{2})$ $=\int\!\mathrm{d}\ell^{-}\,\mathcal{S}\Bigl{(}b_{T},m,\mu,\frac{\ell^{-}}{\nu}\Bigr{)}\,\sqrt{S}(b_{T},m,\mu,\nu)$
@@
where $D^{\mathrm{thr}}_{h/q}$ is the collinear (threshold) fragmentation function, whose renormalization is given by the timelike DGLAP kernels expanded to leading power in $1-z$. The product $C_{m}\,F_{H}$ in eq. (6) is renormalized like a threshold FF [6], thus eq. (6) is indeed RGE consistent by comparison with eq. (6). We note that both eqs. (6) and (6) generically feature a dependence on the heavy quark mass in the soft and collinear-soft function since $mb_{T}\sim 1$, to which we return towards the end of this section.
To make contact with the partonic heavy-quark TMD FF, we next take $\Lambda_{\mathrm{QCD}}\to 0$, which in particular means that now $\Lambda_{\mathrm{QCD}}/m\ll\bar{z}_{H}$. The mode setup for this regime is illustrated in figure -1075 (b). The only change compared to figure -1075 (c) is that the fragmentation shape function now factorizes into the perturbative bHQET shape function $\hat{F}$ and the total fragmentation probability $\chi_{H}$ [6],
$F_{H}(k^{-},\mu)=\hat{F}(k^{-},\mu)\,\chi_{H}+\mathcal{O}(\Lambda_{\mathrm{QCD}})\,.$ (6.5)
@@
$d_{1\,Q/Q}(z_{H},b_{T},\mu,\omega^{2})$ $=C_{m}(m,\mu,\omega^{2})\,\int\!\mathrm{d}\ell^{-}\,\mathcal{S}\Bigl{(}b_{T},m,\mu,\frac{\ell^{-}}{\nu}\Bigr{)}\,\sqrt{S}(b_{T},m,\mu,\nu)$
×$\quad\times\int\!\frac{\mathrm{d}k^{-}}{\bar{n}\cdot v}\,\hat{F}\Bigl{(}\frac{k^{-}}{\bar{n}\cdot v},\mu\Bigr{)}\,\delta\Bigl{(}\bar{z}_{H}-\frac{\ell^{-}+k^{-}}{\omega}\Bigr{)}+\mathcal{O}\bigl{[}(1-z_{H})^{0}\bigr{]}\,.$ (6.6)
This is precisely the relation governing the leading $1/\bar{z}$ terms in the partonic heavy-quark TMD FF that we were seeking out, and it is straightforward to verify eq. (6) at one-loop order against eq. (3.4) using the fact that ~$\tilde{\mathcal{R}}(z,y)=\mathcal{O}[(1-z)^{0}]$, the fact that $\bar{n}\cdot v=\omega/m$, and the ingredients given in appendix B. In particular, since the soft and soft-collinear function are defined as Wilson line correlators, they can only acquire mass dependence through secondary mass effects starting at $\mathcal{O}(\alpha_{s}^{2}C_{F}T_{F})$. In turn, this implies that the mass dependence at $\mathcal{O}(\alpha_{s})$ arises only from that of $C_{m}$, which is logarithmic to all orders, explaining why eq. (3.4) has to have the simple logarithmic mass dependence that we found for $\bar{z}\to 1$. We stress that starting at $\mathcal{O}(\alpha_{s}^{2})$, the soft and soft-collinear functions do develop a highly nontrivial dependence on $mb_{T}$, which is e.g. inherited by the coefficients of $\mathcal{L}_{0}(\bar{z})$ and $\delta(\bar{z})$ in $d_{1\,Q/Q}^{(2)}$ at two loops. The required ingredients are readily available in the literature [44], [^4] The secondary mass effects in $\sqrt{S}$ have been calculated in ref. [44]. For the collinear-soft function one may use that for the exponential regulator of ref. [67], the collinear-soft function is simply given by an exponential of the Collins-Soper kernel to all orders [69]. This remains true in the presence of mass effects in the CS kernel, which have likewise been calculated in ref. [44]. One potential subtlety is that the results of ref. [44] for $\sqrt{S}$ were obtained using the $\eta$ regulator of refs. [39, 40] instead. However, it is easy to see that the one-loop soft function with a massive gluon (which underlies the dispersive calculation of the secondary quark mass effects in ref. [44]) leads to identical results for both regulators, meaning the final renormalized result for the mass-dependent soft function in ref. [44] is also valid for either regulator. which can provide a strong cross check on a future two-loop calculation of the partonic heavy-quark TMD FF, or simplify it by allowing to restrict to finite $z<1$ from the start.

### v3

@@ 4.1 All-order renormalization properties
where $S^{(n_{\ell})}$ is the bare TMD soft function in the theory where only the $n_{\ell}$ light quark degrees of freedom are active. This is the same soft function as appears in light-quark TMD PDFs or FFs at transverse momenta below the heavy-quark threshold: The fact that it also arises here and cancels the rapidity divergences as $\eta\to 0$ in the bare light-spin density matrix follows from consistency of the TMD factorization for flavor-changing hard processes, e.g. charged-current DIS $e^{-}\bar{s}\to\nu_{e}\,\bar{c}$.111 We assume as in section 2.1 that a rapidity regulator $\eta$ is used for which soft (zero-bin) subtractions in the bare collinear matrix element are scaleless. If this is not the case, it is crucial that zero-bin subtractions are defined by modes that are soft in a frame where the heavy quark is highly boosted, i.e., their interactions with the heavy-quark sector are encoded in lightlike Wilson lines and the multiplicative zero-bin subtraction factor is obtained by taking $\bar{n}\cdot v\to\infty$. The UV $\overline{\text{MS}}$ renormalization factor $Z_{\chi_{1}}$ likewise follows from consistency and is given by
@@
A natural boundary condition for $\rho_{\ell}$, where it is free of large rapidity logarithms, is given by $\rho\sim 1$, i.e., the rest frame of the heavy quark. By contrast, in physical cross sections, the boost is fixed to $\rho=Q/m\gg 1$, i.e., the lab frame in the $e^{+}e^{-}$ case, or the Breit frame for SIDIS. The $\rho$ evolution between those scales resums the associated large rapidity logarithms. As in the relativistic case, the two-dimensional evolution equations satisfy an integrability condition,
@@ 4.3 One-loop calculation
Figure -1196: Left: Diagrams contributing to the total bHQET TMD fragmentation factor in the perturbative domain. Wilson lines along the lightlike ${\bar{n}}$ direction are indicated by the slanted double lines. The horizontal double lines indicate the Wilson lines along the timelike $v$ direction. The dashed line indicates on-shell cuts after introducing a complete set of states. The mirror diagram for (a) is understood and included in expressions given in the text. Right: Phase-space integrands $\mathcal{A}$ for the two real-emissions diagrams as a function of the gluon emission rapidity $y$.
@@
in analogy to eq. (3.10). Virtual diagrams with $X$ the vacuum state are scaleless. The two contributing real emission diagrams at $\mathcal{O}(\alpha_{s})$ in Feynman gauge are shown in figure -1196b. The one-gluon Feynman rule for the $Y_{v}$ Wilson line can e.g. be read off from ref. [46], which uses a covariant derivative convention consistent with the one-gluon Feynman rule we use for $W_{\eta}$. Including the mirror diagram for diagram (a), the two diagrams are both given by simple integrals over the rapidity $y$ of the emitted gluon in the heavy-quark rest frame,
@@
The origin of the rapidity divergence, like the structure of the integrands as a whole, has a rather intuitive interpretation, which we illustrate in the right panel of figure -1196b: Emissions at large positive gluon rapidity $y$ in the rest frame of the heavy quark are strongly suppressed for either contributing diagram, which is the well-known dead-cone effect [32]. For diagram (b), where the static heavy quark $Y_{v}$ both emits and absorbs the gluon, the contribution to the emission probability (dot-dashed blue) peaks at $y=0$ and is suppressed also at large negative rapidities because the contribution must be even in the rest frame. In contrast to this, the contribution from diagram (a), where the emission couples to the $W_{\eta}$ that arises from integrating out the interactions with the anticollinear and soft sectors and is required for gauge invariance, tends to a large constant as $y\to-\infty$ (dashed red), leading to a rapidity divergence. Of course, this precisely corresponds to the central soft momentum region, and thus after introducing the regulator $e^{-\eta y}$ with $\eta<0$ to dampen the behavior as $y\to\infty$ (solid red), the associated pole cancels against the soft function as usual.
@@ 6 Joint threshold and $q_{T}$ resummation for massive quark fragmentation
Figure -1195: Relevant modes contributing to the heavy-quark TMD FF as the collinear momentum fraction $z_{H}\to 1$ while counting $m\sim k_{T}$. See the text for a detailed discussion.
@@
We stress that we continue to count $m\sim k_{T}\sim 1/b_{T}$. In figure -1195 we illustrate the contributing modes in the EFT for this scenario. Figure -1195 (a) shows the original, unmodified mode setup that led to eq. (2.1). Here the partonic heavy-quark TMD FF arises from matching collinear modes (green), accounting for the usual contribution of the central soft modes (orange), onto the nonperturbative modes (gray) describing the formation of the brown muck, which is isotropic in the rest frame of the heavy quark indicated by the dashed black line. The collinear modes are kinematically unconstrained in the longitudinal direction because $\bar{z}_{H}\sim 1$.
Once the modified assumption eq. (6.1) is made, energetic collinear modes at the scale $\mu\sim k_{T}\sim m$ can no longer be radiated into the final state, see figure -1195 (b) and (c). Their effect becomes purely virtual and is encoded in a matching coefficient $H_{m,n}^{Q\to h_{v}}$ (green cross) from matching onto bHQET modes at lower virtuality. Real emissions into the $n$-collinear direction at the scale $\mu\sim k_{T}\sim m$ can instead only have collinear-soft (pink) scaling
@@
which is the characteristic mode of double-differential $z\to 1$ threshold and $q_{T}$ resummation, often simply dubbed joint resummation [64, 65, 66, 67, 68, 69, 70, 71, 72].222 We note that while ref. [68] also dealt with final-state heavy quarks, that work considered the joint resummation for the total transverse momentum of a heavy quark pair with respect to the beam direction, which is a distinct problem from the one we consider here.
The physical picture at scales $\mu\ll k_{T}\sim m$ in this scenario, which is governed by one or several bHQET modes at their characteristic lab-frame rapidity $p^{-}/p^{+}\sim Q^{2}/m^{2}$, further depends on the relative size of $\Lambda_{\mathrm{QCD}}/m$ and $\bar{z}_{H}$, specifically whether $\Lambda_{\mathrm{QCD}}/m\sim\bar{z}_{H}$ or $\Lambda_{\mathrm{QCD}}/m\ll\bar{z}_{H}$. (Since $\Lambda_{\mathrm{QCD}}$ introduces a fourth scale into the problem, i.e., a third independent dimensionless power counting parameter, either case may hold in general and independent of $k_{T}/Q$ and $m/Q$.) The relevant modes for the case $\Lambda_{\mathrm{QCD}}/m\sim\bar{z}_{H}$ are illustrated in figure -1195 (c), and are in fact analogous to the well-understood case of single-inclusive heavy-hadron production near the endpoint in $e^{+}e^{-}$ collisions [5, 6]. Specifically, the nonperturbative fragmentation process at the scale $\mu\sim\Lambda_{\mathrm{QCD}}\sim\bar{z}_{H}m$ provides a leading contribution to the measured longitudinal hadron momentum in this case (but not to its transverse momentum, where the contribution is power suppressed!). This effect is encoded in the well-known collinear bHQET fragmentation shape function $F_{H}(k^{-}/v^{-},\mu)$ [5, 6]. [^3] We use the symbol $F_{H}$ here to distinguish the shape function from the TMD soft function; in ref. [6], the symbol $S_{H/Q}$ is used instead, and our partonic $\hat{F}$ corresponds to their $S_{Q/Q}$. In total, the factorization of the hadron-level heavy-quark TMD FF in this regime reads, with $\zeta=\omega^{2}$ the Collins-Soper scale,
$D_{1\,H/Q}(z_{H},b_{T},\mu,\omega^{2})$ $=H_{m,n}^{Q\to h_{v}}\Bigl{(}m,\mu,\frac{\nu}{\omega}\Bigr{)}\,\int\!\mathrm{d}\ell^{-}\,\mathcal{S}(\ell^{-},b_{T},m,\mu,\nu)\,\sqrt{S}(b_{T},m,\mu,\nu)$
×$\quad\times\int\!\frac{\mathrm{d}k^{-}}{v^{-}}\,F_{H}\Bigl{(}\frac{k^{-}}{v^{-}},\mu\Bigr{)}\,\delta\Bigl{(}\bar{z}_{H}-\frac{\ell^{-}+k^{-}}{\omega}\Bigr{)}$
@@
where $\mathcal{S}$ and $\sqrt{S}$ are the separately rapidity-renormalized collinear-soft function from joint resummation [72] and the square root of the standard TMD soft function, respectively. Similarly, $H_{m,n}^{Q\to h_{v}}$ (defined in ref. [47] without the superscript, which we here include for clarity) is the rapidity-renormalized matching coefficient encoding the virtual contributions of massive collinear quark modes $Q$ with label momentum $\omega$ matched onto bHQET fields $h_{v}$. As in ref. [47], the effect of these modes is purely virtual due to the kinematic constraint $z_{H}\to 1$. We note that $H_{m,n}^{Q\to h_{v}}$ also includes the effect of secondary heavy quark loops not directly connected to the operator starting at $\mathcal{O}(\alpha_{s}^{2})$, which are responsible for its rapidity renormalization. By evolving all terms in eq. (6) from their canonical scales (e.g. in Mellin space $N\leftrightarrow z_{H}$, where $\nu_{\mathcal{S}}\sim\mu_{F_{H}}v^{-}\sim\omega/N$ and the evolution is multiplicative) to the common scale $\mu\sim\nu\sim Q$ with the hard function, all large threshold and transverse momentum logarithms present in the heavy-hadron TMD cross section in this regime can be resummed simultaneously.
It is instructive to compare the above to the refactorization of the light-parton TMD FF in the joint resummation regime $k_{T}\ll\bar{z}_{h}Q\ll Q$, which follows from crossing the light-parton TMD PDF refactorization relations [72], accounting for the additional effect of secondary heavy quarks starting at two loops,
$D_{1\,h/q}(z_{h},b_{T},\mu,\omega^{2})$ $=H_{m,n}^{q\to q}\Bigl{(}m,\mu,\frac{\nu}{\omega}\Bigr{)}\int\!\mathrm{d}\ell^{-}\,\mathcal{S}(\ell^{-},b_{T},m,\mu,\nu)\,\sqrt{S}(b_{T},m,\mu,\nu)$
@@
[…]
To make contact with the partonic heavy-quark TMD FF, we next take $\Lambda_{\mathrm{QCD}}\to 0$, which in particular means that now $\Lambda_{\mathrm{QCD}}/m\ll\bar{z}_{H}$. The mode setup for this regime is illustrated in figure -1195 (b). The only change compared to figure -1195 (c) is that the fragmentation shape function now factorizes into the perturbative bHQET shape function $\hat{F}$ and the total fragmentation probability $\chi_{H}$ [6],
$F_{H}\Bigl{(}\frac{k^{-}}{v^{-}},\mu\Bigr{)}=\hat{F}\Bigl{(}\frac{k^{-}}{v^{-}},\mu\Bigr{)}\,\chi_{H}+\mathcal{O}(\Lambda_{\mathrm{QCD}})\,.$ (6.5)
@@
$d_{1\,Q/Q}(z,b_{T},\mu,\omega^{2})$ $=H_{m,n}^{Q\to h_{v}}\Bigl{(}m,\mu,\frac{\nu}{\omega}\Bigr{)}\,\int\!\mathrm{d}\ell^{-}\,\mathcal{S}(\ell^{-},b_{T},m,\mu,\nu)\,\sqrt{S}(b_{T},m,\mu,\nu)$
×$\quad\times\int\!\frac{\mathrm{d}k^{-}}{v^{-}}\,\hat{F}\Bigl{(}\frac{k^{-}}{v^{-}},\mu\Bigr{)}\,\delta\Bigl{(}\bar{z}-\frac{\ell^{-}+k^{-}}{\omega}\Bigr{)}+\mathcal{O}\bigl{[}(1-z)^{0}\bigr{]}\,.$ (6.6)
This is precisely the relation governing the leading $1/\bar{z}$ terms in the partonic heavy-quark TMD FF that we were seeking out, and it is straightforward to verify eq. (6) at one-loop order against eq. (3.4) using the fact that ~$\tilde{\mathcal{R}}(z,y)=\mathcal{O}[(1-z)^{0}]$, the fact that $v^{-}=\omega/m$, and the ingredients given in appendix B. In particular, since the soft and collinear-soft function are defined as Wilson line correlators, they can only acquire mass dependence through secondary mass effects starting at $\mathcal{O}(\alpha_{s}^{2}C_{F}T_{F})$. In turn, this implies that the mass dependence at $\mathcal{O}(\alpha_{s})$ arises only from that of $C_{m}$, which is logarithmic to all orders, explaining why eq. (3.4) has to have the simple logarithmic mass dependence that we found for $\bar{z}\to 1$. We stress that starting at $\mathcal{O}(\alpha_{s}^{2})$, the soft and collinear-soft functions do develop a highly nontrivial dependence on $mb_{T}$, which is thus inherited by the coefficients of $\mathcal{L}_{0}(\bar{z})$ and $\delta(\bar{z})$ in $d_{1\,Q/Q}^{(2)}$ at two loops. The required ingredients are readily available in the literature [74, 48], [^5] The secondary mass effects in $\sqrt{S}$ have been calculated in ref. [48]. For the collinear-soft function one may use that for the exponential regulator of ref. [71], the collinear-soft function is simply given by an exponential of the Collins-Soper kernel to all orders [75]. This remains true in the presence of mass effects in the CS kernel, which have likewise been calculated in ref. [48]. One potential subtlety is that the results of ref. [48] for $\sqrt{S}$ were obtained using the $\eta$ regulator of refs. [43, 44] instead. However, it is easy to see that the one-loop soft function with a massive gluon (which underlies the dispersive calculation of the secondary quark mass effects in ref. [48]) leads to identical results for both regulators, meaning the final renormalized result for the mass-dependent soft function in ref. [48] is also valid for either regulator. which can provide a strong cross check on a future two-loop calculation of the partonic heavy-quark TMD FF, or simplify it by allowing to restrict to finite $z<1$ from the start.